How do you find the domain of f(x)=9/(2+log(3x))?
Answer by stanbon(75887)

find the domain of f(x)=9/(2+log(3x))?
----
Two conditions:
1. The denominator cannot be zero.
2. 3x must be >0
-----------------------------------
1. Solve 2+log(3x) = 0
log(3x) = -2
3x = 10^-2 = 1/100
x = 1/300
-----------
2. 3x > 0
x > 0
-------------
Domain: All Real Numbers >0 and not equal to 1/300.
=====================================================
Cheers,
Stan H.